OpenStudy (anonymous):

Solve the equation: |3x + 15| = 6

OpenStudy (anonymous):

3x = 6 - 15 = -8 x = -8/3

OpenStudy (anonymous):

Sorry, -9/3.

OpenStudy (vishal_kothari):

-7 and -3

OpenStudy (anonymous):

A. x = -3, -7 B. x = 3, -7 C. x = 3, 7 D. x = -3, 7

OpenStudy (anonymous):

How two values?

OpenStudy (vishal_kothari):

a.

OpenStudy (vishal_kothari):

A.

OpenStudy (anonymous):

subtract 15 from both sides you are left with 3x=-9 then divide both sides by three now x=-3

OpenStudy (vishal_kothari):

@pratu square both the sides....

OpenStudy (vishal_kothari):

you will get two values...

OpenStudy (anonymous):

lol thanks

myininaya (myininaya):

3x+15=6 or 3x+15=-6 3x=6-15 3x=-6-15 3x=-9 3x=-21 x=-3 x=-7